OpenStudy (anonymous):

find all the solutions of sec theta - 2 =0. It says use "or" as necessary and gives an example of kpi. I have no idea how to begin this..

OpenStudy (anonymous):

let theta=x so, sec x -2 = 0 sec x = 2 cos x = 1/2 cos x = cos pi/3 x = pi/3 i.e theta = pi/3

OpenStudy (anonymous):

theres other answer too like you said

OpenStudy (anonymous):

Im thinking this, isolate the theta. \[\sec( \theta ) - 2 = 0, subtract 2,\sec ( \theta ) = -2, Divide by secant. ( \theta ) = inverse \sec(-2)\]

OpenStudy (anonymous):

Oh alright, what are the other answers, I got the pi/3

OpenStudy (anonymous):

legoauto is right as is my approach, They both get you to the same result pi/3. its just a matter of preference. I like dividing by whatever function im using to create the inverse of that. the tricky part is just remembering the domains

OpenStudy (anonymous):

killbot, there is another answer too, but idk how to get it.

OpenStudy (anonymous):

This is what the question is exactly asking.

OpenStudy (precal):

another way to look at this is to graph the function

OpenStudy (precal):